OpenStudy (anonymous):

Please someone help me how to solve this. Simplify. 3√(5) -6√(125) Thx!!

OpenStudy (anonymous):

-27√(5)

OpenStudy (anonymous):

Hi, Fontez. Do you have the step?? Thx!!

OpenStudy (anonymous):

125/5=25 25/5=5 5/5=1 ok? √(125) = 5√(5) 3√(5) -6*5√(5) 3√(5) -30√(5) 27√(5)

OpenStudy (anonymous):

resolves?

OpenStudy (anonymous):

Thx so much!!

OpenStudy (anonymous):

sorry fault sinal :) -27√(5)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Thx Now I understand. :D